OpenStudy (anonymous):

cos theta = 4/5, 0˚< theta < 90˚ use the information given to find sin2theta, cos2theta, and tan 2theta

OpenStudy (anonymous):

you need \(\sin(\theta)\) so compute these do you know how to find it?

OpenStudy (anonymous):

the identities?

OpenStudy (anonymous):



OpenStudy (anonymous):

sin 2 theta = 2 sin theta cos theta cos 2 theta = 2 cos ^2 theta - 1 tan 2 theta = sin 2 theta / cos 2 theta

OpenStudy (anonymous):

do i plug in 4/5 where theta is?

OpenStudy (anonymous):

If costheta=4/5 then sintheta= 3/5 \[\sin2\theta=2\sin\theta \cos\theta\]

OpenStudy (anonymous):

there is a picture of an angle whose cosine is \(\frac{4}{5}\) you find the third side by pythagoras, or recalling the 3 - 4 - 5 right triangle this tells you \(\sin(\theta)=\frac{3}{5}\)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so 2 sin theta = 2 sin 3/5 cos 4/5

OpenStudy (anonymous):

no do not make that mistake, it is a common one do not use \(\theta=\frac{4}{5}\) use \(\cos(\theta)=\frac{4}{5}\) so \[\sin(2\theta)=2\times \frac{4}{5}\times \frac{3}{5}\]

OpenStudy (anonymous):

no not the cosine and sine of those numbers, those numbers ARE the cosine and sine

OpenStudy (anonymous):

use the numbers themselves, they are your sines and cosines

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

your good now right?

OpenStudy (anonymous):

im still confused a little ..